You are here

Beam delivery update

header

05 December 2014

The BL1A current had been ~125 µA for the last several days to ensure a decent neutron flux at TNF for outside users. That’s no longer needed, so the BL1A current has been reduced to ~110 µA. BL2A is running at 9.8 µA while BL2C1 continues to take nA as needed for PIF; this is all expected to continue through the weekend.

At ISAC, the mass separator room was accessed earlier today in preparation for next week’s planned RIB delivery; 47K delivery to GRIFFIN resumed just after 16:00 and will continue until 08:30 Monday. Tuning to the ISAC implantation station for experiment L112 will begin at that time. Meanwhile, 76Se has been delivered to DRAGON from the OLIS multicharge ion source for measurements of Se charge-state distributions; those are expected to continue through the weekend, but only during extended daytime hours – beam won’t be delivered overnight as there’s no need for it.

-- ACM